Mr. Coleman, an unmarried individual, has the following income items: Interest income $ 24,700 Ordinary loss from an S corporation (8,650 ) Ordinary income from a partnership 64,000 He has $10,250 itemized deductions and no dependents. Mr. Coleman's passthrough income is qualified business income under Section 199A. Compute Mr. Coleman’s income tax. Assume the taxable year is 2019. Use Individual Tax Rate Shedules and Standard Deduction Table. (Round your answers to the nearest whole dollar amount.)
Answer #1

AGI = interest income - ordinary loss from an S corporation + ordinary income from a partnership = 24700 - 8650+64000 = 80050

Income limit for pass-through businesses for being eligible for section 199A deduction is $160,700 to $210,700 in case of single filers for the year 2019. As the income is below $160700 so no such deduction.

Taxable income = AGI – (greater of standard deduction or itemized deduction, 12200 or 10250) = 80050-12200 = 67850

Income tax liability (using tax rates for single) = 4543+(22%*(67850-39475)) = 10785.50 = $10786